OpenStudy (anonymous):

find the discriminate of -4x^2-3x+5=0

Directrix (directrix):

How to find the discriminant. See attachment.

OpenStudy (anonymous):

i know but i got -71

OpenStudy (isaiah.feynman):

Do you think its wrong?

OpenStudy (anonymous):

89

OpenStudy (anonymous):

yea because its not one of the aswers a)90 b)91 c)\[\sqrt{80}\] d89

OpenStudy (anonymous):

how did u get that @FutureMathProfessor

OpenStudy (nurali):

-4x^2-3x+5=0 a = -4 b = -3 c = 5 Discriminant: b^2-4ac b^2-4ac = (-3)^2-4*(-4)*5 = 9+80 = ?